OpenStudy (anonymous):

y varies inversely with twice the value of x. When y = 10, x = 3. Find y when x = 5.

OpenStudy (anonymous):

y=k/(2x) find k using y=10 and x=3

OpenStudy (anonymous):

12 , 4 ,10 ,6?

OpenStudy (anonymous):

im guessing those r the answer choices do u want help or the answer?

OpenStudy (anonymous):

answer because its timed

OpenStudy (anonymous):

6

OpenStudy (anonymous):

thankssss

OpenStudy (anonymous):

y=k/(2x) 10=k/(2*3) 10*2*3=k k=60 y=60/(2x)=30/x y=30/5=6

OpenStudy (anonymous):

np